OpenStudy (anonymous):

find y' and y'' y=cos^2x I have y' as 2(cosx)(-sinx)=-sin2x for y'' all I can come up with is -cos2x-sin2, is that right?

OpenStudy (turingtest):

y' is correct for y'' you failed to use the chain rule let u=2x-> u'=? \[(-\sin(u))'=-\cos(u)\cdot u'\]

OpenStudy (anonymous):

ah, got it, I used the product rule instead of the chain rule

OpenStudy (anonymous):

thank you

OpenStudy (turingtest):

welcome